# Empowering Agile-Based Generative Software Development through Human-AI Teamwork

### v1

@@ 1. Introduction
Figure 1. Problem Statement Diagram. Users are unsure how to drive the “Agent” to generate desired software. The “Agent” does not know how to fulfill user requirements. We have built a bridge between users and the “Agent”, facilitating collaboration between human decision-making skills and the Agent’s coding capabilities. This collaboration has created a generative software development “Agent” with lightweight iterative feedback.
@@ 3. Design of Human-AI Collaborative AgileGen
In this section, we first outline the architectural design of AgileGen, an Agile-based generative software development through human-AI teamwork. Then, in Section 3.2, we will introduce the end-user decision-making and interaction interface components in AgileGen. Finally, in Section 3.3, we will detail the design of the core components of AgileGen. We use website application development as an example to demonstrate how AgileGen navigates end-user decisions through various components to generate a website application that aligns with user requirements.
Additionally, this section provides an example to illustrate how AgileGen bridges collaborative development between end users and large language models, aiming to fulfill a single-line requirement: ”Please generate a web system with random roll call function.”
@@ 3.1. AgileGen Overview
(6) The decided natural language scenarios are converted back into Gherkin language ($Gherkin^{*}$) and received by the next component.
Then, the decided $Gherkin^{*}$ is received by the Rapid Prototype Design component, with the workflow as follows:
(7) Plan visual design guided by $Gherkin^{*}$.
(8) Generate consistency factors consistent with the requirements guided by $Gherkin^{*}$.
(9) Integrate visual design, and $Gherkin^{*}$ to guide large language models in generating code.
@@ 3.2.1. User Decision-Making
Figure 6. AgileGen Interaction Interface that Aids End-User Decision-Making.
@@ 3.3.1. Scenario Design component
The Scenario Design component is used to supplement the missing acceptance criteria with the end-user to clarify the end-user’s requirements. The detailed implementation is shown in Figure 7. Here, 1, 2, 3, and 4 represent different functional blocks within the component, with the dashed line enclosing the output of these blocks. After the end-user submits a requirement $nl$, the Scenario Design component follows the sequence of End User requirement decision-making $\rightarrow$ 1: Memory Pool $\rightarrow$ 2: Scenario Design $\rightarrow$ 3: Gherkin to NL $\rightarrow$ End User Scenarios decision-making $\rightarrow$ 4: NL to Gherkin, ultimately outputting the decided $Gherkin^{*}$.
@@
The memory pool is initially empty and gets populated with data as users make scenario decisions. After a user makes a decision, the user’s requirement $nl$ along with the decided scenarios $NL^{*}=\{nl^{*}(S_{1}),nl^{*}(S_{2})\}$ are recorded in the memory pool as human experiential knowledge. While there may be multiple real scenarios, $S_{1}$ and $S_{2}$ are used as examples for convenience. The continuously accumulated human experiential knowledge serves as examples for generating subsequent scenarios. By designing such a memory pool, we aim to enhance the reliability of our system’s output through ongoing practice.
@@
2: Scenarios Design Agile utilizes user story to acquire demands for software development. These stories require users to specify their role design, functionality vision, and value statement. Acceptance criteria are important for determining whether the user expectations are met before project delivery (Dees et al., 2013). However, not everyone is good at storytelling. To address this, Gherkin uses keywords to structure and define executable specifications for the user story (Parsa, 2023) and helps define acceptance criteria for end-user requirements. The goal is to generate a well-formed Gherkin language from the $nl$ of the end-user. The entire scenario design AI-Chain is depicted in Figure 7(2), including selected scenario examples from the memory pool, a predefined Gherkin template (Parsa, 2023) (Feature:[Title]), a Gherkin instance (Feature: Adding-Removing-items), and the input interface Feature:{$nl$}. Moreover, explanatory comments are added to the Gherkin template to guide the large language models in better understanding and generating the user story in Gherkin format corresponding to $nl$, following the Chain-of-thought concept (Wei et al., 2022). The generated result includes two scenarios, as shown in the Gherkin dashed expansion box in Figure 7. The user story clearly describes the motivation, action, and feedback, with well-defined acceptance criteria.
@@
3,4: Interactive Bridge with Gherkin The user story, written in formal Gherkin language, describes the functionalities needed. During the scenarios decision-making, communication with the end-user for clarifications requirements. End-users tend to be more at ease when communicating in non-formal natural language (NL). This section explains how we will design an interaction bridge to eliminate the gap between non-formal natural language and formal Gherkin.
@@
3: Gherkin to NL The process of converting the Gherkin language into natural language descriptions is illustrated in Figure 7(3). Gherkin from 2: Scenario Design is segmented by rules introduced in Section 3.2.2. The segmented scenarios are filled into the input interface of 3: Gherkin to NL as Secnarios:[$S_{1},S_{2}$]. This is followed by an instruction we found effective through multiple trials: generating natural language from Gherkin and attempting to explain it. Our agent interpret the “Scenarios” in this instruction corresponding to the contents after the keyword ”Scenarios: ”. The generated results are displayed in the NL dashed expansion box in Figure 7, with natural language scenarios arranged by sequence number.
@@
4: NL to Gherkin The process of converting the natural language scenarios decided by the user, $NL^{*}=<nl^{*}\left(S_{1}\right),nl^{*}\left(S_{2}\right)>$, into Gherkin scenarios is depicted in Figure 7(4), consisting of four parts: the natural language input interface, Instruction, Gherkin template, and Gherkin instance. Unlike the “Scenarios Design”, the instruction here focuses on converting natural language into Gherkin scenarios, not the entire feature. The user-decided scenarios $NL^{*}$ are input into the “4: NL to Gherkin” input interface, and recorded in the memory pool along with the end-user’s requirement $nl$. For example, as shown in Figure 7, the end-user modified Scenario 1 during the scenario decision-making, adding the sentence: ”Then the chosen student could be any of the five students.” The decided scenario $NL^{*}$ and the end-user requirement $nl$ are recorded as $Item_{18}$ and saved in the memory pool. Then, the decided scenarios $NL^{*}$ are input into the 4: NL to Gherkin natural language input interface and the final Gherkin scenarios are represented by $Gherkin^{*}$. With this, the Scenario Design component completes its task.
3.3.2. Rapid Prototype Design component
@@
The Rapid Prototype Design component is tasked with generating software prototypes from Gherkin. The implementation details, as shown in Figure 8, follow the output $Gherkin^{*}$ from the above Scenario Design component, proceeding through 5: Visual Design $\rightarrow$ 6: Code Generation $\rightarrow$ 7: Consistency Factor $\rightarrow$ 8: Auto Modification $\rightarrow$ End User Acceptance and Recommendation Decision-making $\rightarrow$ 9.1: Design Modification / 9.2: Code Modification, to ultimately output the software’s code. Here, 9.1 and 9.2 represent two interactive choices for modifications, namely visual design modifications and functional code modifications, respectively. After the code is executed and presented to the end-user, they can choose and input the appropriate modification method based on their needs. The modified code will be executed and presented again for the end-user to make decisions. Notably, our method interacts with the end-user iteratively until the user accepts the generated result.
@@
To ensure the integrity of the $Gherkin^{*}$ file, we merge the separate scenarios generated by NL to Gherkin and prepend the file with the keyword “Feature:$nl$” at the beginning.
@@
Page design is a key part of the site organization in the research-based Web Design and Usability Guidelines (Leavitt and Shneiderman, 2006). As shown on the left of Figure 8(5), this part designs the page content description for the website application through $Gherkin^{*}$. Specifically, $Gherkin^{*}$ is input into the “Scenario Description:” and an instruction guides the large language models to generate a page design description. The generated result is shown in the $Page~{}design$ dashed expansion box in Figure 8, including page content information such as Header, Form, Footer, etc. Particularly, Sitemaps describe the URL structure of the webpage.
The visual description, as shown on the right of Figure 8(5), takes $Page~{}design$ as input, with instructions and eight principles guiding the large language models to design the visual description. The incorporation of 8 successful visual design principles [^12] https://www.usability.gov/what-and-why/visual-design.html aims to guide the generation of descriptions centered around these principles. The generated result of the visual description is displayed in Figure 8 under Visual description. Our agent have generated a comprehensive visual description based on the page design. It outlines which content and functions should be emphasized.
6: Code Generation. The purpose of code generation is to guide the large language models to generate code based on the Task, which stems from the end-user’s raw requirement description. As shown in Figure 8(6), the AI-Chain design for code generation includes Gherkin and Visual Design input interfaces. These interfaces respectively receive $Gherkin^{*}$ containing human decisions and Page design & Visual description related to visual design. In practice, the large language models’ generation results are diverse, and to facilitate the segmentation of code files, we also prescribe the starting format for generating code. Underneath the instructions, we add ”Please generate the codes for the three files in ¡Code¿ without any note” and the following format (taking a website application as an example): 1.index.html: “‘html ¡Code¿ “‘ end index.html 2.style.css: “‘css ¡Code¿ “‘ end style.css 3.script.js: “‘javascript ¡Code¿ “‘ end script.js In this way, we can utilize the code file extraction rules described in the Acceptance and Recommendation Decision-making section referred to in 3.2.2, to segment different code files. The generated code files are denoted by “Code”.
@@
Summary: In the Scenario Design component, from steps 1 to 4, the end-user is involved in proposing requirements and making scenario decisions. From the end-user raw requirement $nl$, a requirement with acceptance criteria $Gherkin^{*}$ is obtained. $Gherkin^{*}$ is used as the input for the Rapid Prototype Design component, where steps 5 to 8 involve automatic rapid prototype designing, executed automatically by AI without the need for end-user participation. Steps 9.1 and 9.2 are selected based on modification recommendations proposed by the end-user, meaning the user triggers different chains by choosing different textboxes for input and clicking different buttons on the interaction interface. Therefore, we adhere to the principle of positioning the end-users involvement at both ends of the process, with AI participation set in the middle. This approach fully leverages the strengths of both end-users and AI, collaborating in a Jigsaw puzzle-like manner to complete generative software development.
@@ 5.1. RQ1:How effective are the software applications generated by our framework?
  • For ”NewsMeter,”  [^21] https://github.com/HarrisClover/AgileGen-SRDD-Experiment/tree/main/NewsMeter the task description is ”Evaluate the credibility of news articles by analyzing multiple factors and generating trustworthiness scores with explanations and evidence.” Our method creates an application that allows users to upload articles and see evaluation results after clicking the ”Evaluate Article” button. ChatDev fails to produce an effective interface. We found that the displayed analysis results are a simulated process, which we believe is reasonable, as evaluating the credibility of news is a complex function that may require integrating suitable APIs and providing more detailed information. Our method generates a software application that simulates the necessary functional processes, laying an important foundation for developing web applications capable of implementing complex features.
  • The task description for ”VideoClipper”  [^22] https://github.com/HarrisClover/AgileGen-SRDD-Experiment/tree/main/VideoClipper is ”VideoClipper is a software application that allows users to easily clip and trim videos. It provides an intuitive interface to select specific sections of the video and saves the trimmed video as a new file.” All methods generate projects capable of video clipping, but our method demonstrates better interface design and interactivity.
@@ 5.3. RQ3: What is the human-computer interaction experience for end-users during the AgileGen process?
Figure 18. The interaction showcases for the ”09 Sound Board” project includes AgileGen’s interaction system on the left and the generated end-user web application interface on the right. The participant made five iterative modifications to the project, including visual design and functionality changes.
@@ 6. Discussion
1) Potential Benefits of Integrating External Tools into Our Method: Our method integrates the SQLite3 external database tool in the memory pool to save human decision experiences. SQLite3 was chosen because it is an open-source, lightweight, embedded relational database management system that does not require server configuration. SQLite3 supports most SQL standard syntax, allowing us to use SQL syntax to access human experiences in the database. Of course, if we need high concurrency and complex query functions in the future, it can easily be integrated with server-side database management tools. More importantly, our method is based on behavior-driven development (BDD), with the scenario design module containing users’ clarified Gherkin scenarios. These Gherkin scenarios serve as test cases, and in the future, we can directly integrate the external Cucumber testing tool to conduct tests. The feedback from these test cases can be used to improve code quality. Since Gherkin focuses more on the consistency between implementation code and business logic, introducing the Cucumber testing tool will help generate software applications that align with user expectations, reducing manual iterations. For example, the initial scenario generation in the ”09 Sound Board” project includes ”Then the website plays the pronunciation of the word ’Apple.’” However, ”Round 1” does not generate the pronunciation functionality. If we introduce the external Cucumber tool, it would automatically test and modify the functionality, reducing the need for additional manual iterations. This will be a key focus of our future research.
@@
2) “Scenario Decision-making”: This involves users adding, deleting, or modifying the generated Gherkin scenarios in our system to clarify their requirements further. We designed a memory pool module to reduce the time cost for users when modifying scenarios. This module accumulates previous user decision experiences and recommends similar scenarios for future similar projects. However, to ensure fairness in the experimental setup, we did not provide any human experience before testing all projects, so the impact of the memory pool can only be measured later. Our design mechanism accumulates human experiences over time, reducing user decision costs as more tests are conducted. In the future, we plan to deploy our method on a public testing platform to collect more human experiences over time, demonstrating the process of continuous enhancement through human-in-the-loop interactions. Additionally, we built an interaction bridge to reduce the learning cost for users when adding and modifying Gherkin scenarios, allowing users to add and modify functionalities using natural language seamlessly.

### v2

@@ Abstract.
S. Zhang also with CSIRO’s Data61
This work was supported by the Project of Science and Technology Research and Development Plan of China Railway Corporation (N2023J044).
@@ 1. Introduction
Figure 1. Problem Statement Diagram. Users are unsure how to drive the Agent to generate desired software, and the Agent does not know how to fulfill user requirements. We have built a bridge between users and the Agent, facilitating collaboration between human decision-making skills and the Agent’s coding capabilities. This collaboration has created a generative software development Agent with lightweight iterative feedback.
@@ 3. Design of Human-AI Collaborative AgileGen
In this section, we first outline the architectural design of AgileGen, an Agile-based generative software development through human-AI teamwork. Then, in Section 3.2, we will introduce the end-user decision-making and interaction interface components in AgileGen. Finally, in Section 3.3, we will detail the design of the core components of AgileGen. We use website application development as an example to demonstrate how AgileGen navigates end-user decisions through various components to generate a website application that aligns with user requirements. Additionally, this section provides an example to illustrate how AgileGen bridges collaborative development between end users and large language models, aiming to fulfill a single-line requirement: ”Please generate a web system with random roll call function.”
@@ 3.1. AgileGen Overview
(6) The decided natural language scenarios are converted back into Gherkin language ($\mathrm{Gherkin^{*}}$) and received by the next component.
Then, the decided $\mathrm{Gherkin^{*}}$ is received by the Rapid Prototype Design component, with the workflow as follows:
(7) Plan visual design guided by $\mathrm{Gherkin^{*}}$.
(8) Generate consistency factors consistent with the requirements guided by $\mathrm{Gherkin^{*}}$.
(9) Integrate visual design and $\mathrm{Gherkin^{*}}$ to guide large language models in generating code.
@@ 3.2.1. User Decision-Making
Figure 6. AgileGen Interaction Interface that helps End-User Decision-Making.
@@ 3.3.1. Scenario Design component
The Scenario Design component is used to supplement the missing acceptance criteria with the end-user to clarify the end-user’s requirements. The detailed implementation is shown in Figure 7. Here, 1, 2, 3, and 4 represent different functional blocks within the component, with the dashed line enclosing the output of these blocks. After the end-user submits a requirement $nl$, the Scenario Design component follows the sequence of End User requirement decision-making $\rightarrow$ 1: Memory Pool $\rightarrow$ 2: Scenario Design $\rightarrow$ 3: Gherkin to NL $\rightarrow$ End User Scenarios decision-making $\rightarrow$ 4: NL to Gherkin, ultimately outputting the decided $\mathrm{Gherkin^{*}}$.
@@
The memory pool is initially empty and gets populated with data as users make scenario decisions. After a user makes a decision, the user’s requirement $nl$ along with the decided scenarios $\mathrm{NL^{*}}=\{nl^{*}(S_{1}),nl^{*}(S_{2})\}$ are recorded in the memory pool as human experiential knowledge. While there may be multiple real scenarios, $S_{1}$ and $S_{2}$ are used as examples for convenience. The continuously accumulated human experiential knowledge serves as examples for generating subsequent scenarios. By designing such a memory pool, we aim to enhance the reliability of our system’s output through ongoing practice.
@@
2: Scenarios Design. Agile utilizes user story to acquire demands for software development. These stories require users to specify their role design, functionality vision, and value statement. Acceptance criteria are important for determining whether the user expectations are met before project delivery (Dees et al., 2013). However, not everyone is good at storytelling. To address this, Gherkin uses keywords to structure and define executable specifications for the user story (Parsa, 2023) and helps define acceptance criteria for end-user requirements. The goal is to generate a well-formed Gherkin language from the $nl$ of the end-user. The entire scenario design AI-Chain is depicted in Figure 7(2), including selected scenario examples from the memory pool, a predefined Gherkin template (Parsa, 2023) (Feature:[Title]), a Gherkin instance (Feature: Adding-Removing-items), and the input interface Feature:{$nl$}. Moreover, explanatory comments are added to the Gherkin template to guide the large language models in better understanding and generating the user story in Gherkin format corresponding to $nl$, following the Chain-of-thought concept (Wei et al., 2022). The generated result includes two scenarios, as shown in the Gherkin dashed expansion box in Figure 7. The user story clearly describes the motivation, action, and feedback, with well-defined acceptance criteria.
@@
3,4: Interactive Bridge with Gherkin. The user story, written in formal Gherkin language, describes the functionalities needed. During the scenarios decision-making, communication with the end-user for clarifications requirements. End-users tend to be more at ease when communicating in non-formal natural language (NL). This section explains how we will design an interaction bridge to eliminate the gap between non-formal natural language and formal Gherkin.
@@
3: Gherkin to NL. The process of converting the Gherkin language into natural language descriptions is illustrated in Figure 7(3). Gherkin from 2: Scenario Design is segmented by rules introduced in Section 3.2.2. The segmented scenarios are filled into the input interface of 3: Gherkin to NL as Secnarios:[$S_{1},S_{2}$]. This is followed by an instruction we found effective through multiple trials: generating natural language from Gherkin and attempting to explain it. Our agent interpret the “Scenarios” in this instruction corresponding to the contents after the keyword ”Scenarios: ”. The generated results are displayed in the NL dashed expansion box in Figure 7, with natural language scenarios arranged by sequence number.
@@
4: NL to Gherkin. The process of converting the natural language scenarios decided by the user, $NL^{*}=<nl^{*}\left(S_{1}\right),nl^{*}\left(S_{2}\right)>$, into Gherkin scenarios is depicted in Figure 7(4), consisting of four parts: the natural language input interface, Instruction, Gherkin template, and Gherkin instance. Unlike the “Scenarios Design”, the instruction here focuses on converting natural language into Gherkin scenarios, not the entire feature. The user-decided scenarios $\mathrm{NL^{*}}$ are input into the “4: NL to Gherkin” input interface, and recorded in the memory pool along with the end-user’s requirement $nl$. For example, as shown in Figure 7, the end-user modified Scenario 1 during the scenario decision-making, adding the sentence: ”Then the chosen student could be any of the five students.” The decided scenario $\mathrm{NL^{*}}$ and the end-user requirement $nl$ are recorded as $Item_{18}$ and saved in the memory pool. Then, the decided scenarios $\mathrm{NL^{*}}$ are input into the 4: NL to Gherkin natural language input interface and the final Gherkin scenarios are represented by $\mathrm{Gherkin^{*}}$. With this, the Scenario Design component completes its task.
3.3.2. Rapid Prototype Design component.
@@
The Rapid Prototype Design component is tasked with generating software prototypes from Gherkin. The implementation details, as shown in Figure 8, follow the output $\mathrm{Gherkin^{*}}$ from the above Scenario Design component, proceeding through 5: Visual Design $\rightarrow$ 6: Code Generation $\rightarrow$ 7: Consistency Factor $\rightarrow$ 8: Auto Modification $\rightarrow$ End User Acceptance and Recommendation Decision-making $\rightarrow$ 9.1: Design Modification / 9.2: Code Modification, to ultimately output the software’s code. Here, 9.1 and 9.2 represent two interactive choices for modifications, namely visual design modifications and functional code modifications, respectively. After the code is executed and presented to the end-user, they can choose and input the appropriate modification method based on their needs. The modified code will be executed and presented again for the end-user to make decisions. Notably, our method interacts with the end-user iteratively until the user accepts the generated result.
@@
To ensure the integrity of the $\mathrm{Gherkin^{*}}$ file, we merge the separate scenarios generated by NL to Gherkin and prepend the file with the keyword “Feature:$nl$” at the beginning.
@@
Page design is a key part of the site organization in the research-based Web Design and Usability Guidelines (Leavitt and Shneiderman, 2006). As shown on the left of Figure 8(5), this part designs the page content description for the website application through $\mathrm{Gherkin^{*}}$. Specifically, $\mathrm{Gherkin^{*}}$ is input into the “Scenario Description:” and an instruction guides the large language models to generate a page design description. The generated result is shown in the $Page~{}design$ dashed expansion box in Figure 8, including page content information such as Header, Form, Footer, etc. Particularly, Sitemaps describe the URL structure of the webpage. The visual description, as shown on the right of Figure 8(5), takes $Page~{}design$ as input, with instructions and eight principles guiding the large language models to design the visual description. The incorporation of 8 successful visual design principles [^12] https://www.usability.gov/what-and-why/visual-design.html aims to guide the generation of descriptions centered around these principles. The generated result of the visual description is displayed in Figure 8 under Visual description. Our agent have generated a comprehensive visual description based on the page design. It outlines which content and functions should be emphasized.
6: Code Generation. The purpose of code generation is to guide the large language models to generate code based on the Task, which stems from the end-user’s raw requirement description. As shown in Figure 8(6), the AI-Chain design for code generation includes Gherkin and Visual Design input interfaces. These interfaces respectively receive $\mathrm{Gherkin^{*}}$ containing human decisions and Page design & Visual description related to visual design. In practice, the large language models’ generation results are diverse, and to facilitate the segmentation of code files, we also prescribe the starting format for generating code. Underneath the instructions, we add ”Please generate the codes for the three files in ¡Code¿ without any note” and the following format (taking a website application as an example): 1.index.html: “‘html ¡Code¿ “‘ end index.html 2.style.css: “‘css ¡Code¿ “‘ end style.css 3.script.js: “‘javascript ¡Code¿ “‘ end script.js In this way, we can utilize the code file extraction rules described in the Acceptance and Recommendation Decision-making section referred to in 3.2.2, to segment different code files. The generated code files are denoted by “Code”.
@@
Summary: In the Scenario Design component, steps 1 to 4 involve the end-user in proposing requirements and making scenario decisions. From the end-user’s raw requirement $nl$, a requirement with acceptance criteria, $\mathrm{Gherkin^{*}}$, is generated. This $\mathrm{Gherkin^{*}}$ serves as the input for the Rapid Prototype Design component, where steps 5 to 8 involve automatic prototype design by AI without end-user involvement. Steps 9.1 and 9.2 are triggered based on modification recommendations from the end-user, allowing them to choose different inputs and actions on the interface. This approach places end-user involvement at both the beginning and end of the process, with AI handling the middle stages. By doing so, it maximizes the strengths of both end-users and AI, working together like pieces of a jigsaw puzzle to achieve generative software development.
@@ 5.1. RQ1:How effective are the software applications generated by our framework?
  • For ”NewsMeter,”  [^21] https://github.com/HarrisClover/AgileGen-SRDD-Experiment/tree/main/NewsMeter the task is to ”Evaluate the credibility of news articles by analyzing multiple factors and generating trustworthiness scores with explanations and evidence.” Our method creates an application where users can upload articles and view evaluation results by clicking the ”Evaluate Article” button. Unlike ChatDev, which fails to produce an effective interface, our approach simulates the analysis process, which we find reasonable. Evaluating news credibility is complex and may require integrating APIs and providing detailed information. By simulating these functional processes, our method provides a solid foundation for developing web applications with advanced features.
  • For ”VideoClipper”  [^22] https://github.com/HarrisClover/AgileGen-SRDD-Experiment/tree/main/VideoClipper is a software application that allows users to easily clip and trim videos. It provides an intuitive interface to select specific sections of the video and saves the trimmed video as a new file.” All methods generate projects capable of video clipping, but our method demonstrates better interface design and interactivity.
@@ 5.3. RQ3: What is the human-computer interaction experience for end-users during the AgileGen process?
Figure 18. The interaction showcases for the ”09 Sound Board” project includes AgileGen’s interaction system on the left and the generated end-user web application interface on the right. The participant made five iterative modifications to the project.
@@ 6. Discussion
1) Potential Benefits of Integrating External Tools into Our Method: Our method integrates the SQLite3 external database tool in the memory pool to save human decision experiences. We chose SQLite3 because it is an open-source, lightweight, embedded relational database that requires no server configuration and supports most SQL syntax, enabling easy access to stored experiences. If we need higher concurrency or more complex queries in the future, SQLite3 can easily integrate with server-side database management tools. Our approach is based on behavior-driven development (BDD), where the scenario design module includes user-defined Gherkin scenarios as test cases. These scenarios can be tested using the external Cucumber tool, which would provide feedback to improve code quality. Since Gherkin ensures alignment between code and business logic, integrating Cucumber can help create software that meets user expectations, reducing manual iterations. For example, in the ”09 Sound Board” project, the initial scenario specifies, ”Then the website plays the pronunciation of the word ‘Apple.’” However, in ”Round 1,” this functionality is not generated. By introducing Cucumber, the tool could automatically test and correct such functionality, minimizing manual corrections. This will be a key focus of our future research.
@@
2) “Scenario Decision-making”: This involves users adding, deleting, or modifying the generated Gherkin scenarios in our system to clarify their requirements further. We designed a memory pool module to reduce the number of decision iterations caused by poor scenario quality. This module accumulates previous user decision experiences and recommends similar scenarios for future similar projects. However, to ensure fairness in the experimental setup, we did not provide any human experience before testing all projects, so the impact of the memory pool can only be measured later. Our design mechanism accumulates human experiences over time, reducing user decision costs as more tests are conducted. In the future, we plan to deploy our method on a public testing platform to collect more human experiences over time, demonstrating the process of continuous enhancement through human-in-the-loop interactions. Additionally, we built an interaction bridge to reduce the learning cost for users when adding and modifying Gherkin scenarios, allowing users to add and modify functionalities using natural language seamlessly.